# Neural timescales from a computational perspective

### v1

@@ 3 Mechanistic models of timescales in the brain
Mechanistic models, when constrained by experimental observations of neural timescales, can provide insights into how the diversity of timescales arises within and across different brain areas. Such models have so far been used to explain the following experimentally observed properties of neural timescales: (i) the hierarchical structure of neural timescales across different brain areas[46], (ii) the diversity of timescales within one brain area[47, 48, 8], and (iii) the flexibility of timescales to change according to different task variables[8, 49, 50, 51]. Mechanisms for explaining the diversity of neural timescales are generally categorized into cellular and synaptic mechanisms on the level of individual neurons and network-mediated mechanisms through interactions between neurons (Fig. 3). While these mechanisms define the “intrinsic” timescales of neurons, other processes, such as external input and neuromodulatory mechanisms, can dynamically modulate them in addition. Moreover, different models capture neural timescales on different spatial scales of observations, e.g., timescales of single neurons versus population activity. Thus, comparing these models can reveal how neural timescales across different scales and modalities can be related to each other.
Box 1. Timescale as an estimated quantity vs. time constant as a parameter Neural timescales in the literature, for the most part, refer to the estimated decay time constant of measured neural activity, i.e., by describing the autocorrelation function of spike counts or LFP time series as $\textrm{ACF}(t)=e^{\frac{t}{\tau_{\textrm{fit}}}}$. Methods discussed in Section 2 generally aim to find the value of $\tau_{\textrm{fit}}$ (or multiple $\tau_{\textrm{fit}}$s) that best fit the data. On the other hand, mechanistic (i.e. Hodgkin-Huxley) and functional (i.e. rate-based artificial recurrent neural networks) neural models in Sections 3 and 4 typically include a time constant parameter that describes how the impact of input into a neuron decays over time, e.g., $\tau_{\textrm{param}}\frac{dV}{dt}=-(V(t)-V_{\textrm{rest}})+I(t)$, where $I(t)$ includes time-varying stimulus-driven, feed-forward, and recurrent inputs. It is important to note that the former, estimated $\tau_{\textrm{fit}}$ does not necessarily need to correspond to the latter, underlying model parameter $\tau_{\textrm{param}}$, and in general, do not. Section 3 outlines multiple mechanisms, including non-time constant related ones such as network connectivity, that contribute to shaping $\tau_{\textrm{fit}}$ we observe in data.
@@ 3.1 Cellular and synaptic mechanisms
In computational models, timescales are often controlled by incorporating various biophysical time constants of cellular and synaptic processes in the dynamics of individual neurons that reflect heterogeneous properties of ionic channels and synapses[52, 53]. Most dynamical system models include a time constant in the dynamical equation of each neuron, interpreted as the membrane time constant ($5-50$ ms[54]) (Box 1). Slower timescales can be implemented by introducing other biophysical features such as slow synaptic time constants[55, 30], timescales of short-term synaptic depression and facilitation [56, 57], or adaptation mechanisms[58, 59, 60, 30]. Spike-frequency adaptation, in particular, spans a wide range of fast and slow timescales following a power law, which is suggested to lead to efficient coding in individual cortical neurons driven by behaviorally relevant signals[60] and facilitate computations on temporally dispersed inputs [61].
@@
However, biophysical time constants at the single-neuron level cannot explain all the temporal structures of neural dynamics. For instance, a Hodgkin-Huxley model of an isolated neuron with detailed dynamics of membrane potential and ionic and synaptic currents cannot capture the full autocorrelation shape of neural activity recorded in prefrontal areas[62]. Biophysical time constants alone define the autocorrelation timescale of neural activity—henceforth referred to as activity timescale (Box 1)—when a neuron is isolated, but when neurons are embedded in a network, their activity timescale may differ significantly from their biophysical time constants. At the same time, distinct biophysical mechanisms may interact differently in the presence of network interactions compared to isolated neurons. In particular, while activity timescales of network-embedded neurons increase with the synaptic time constant[30], their relationship to the adaptation time constant is less straightforward: Analyzing similar models of randomly connected excitatory and inhibitory units, Beiran and Ostojic[30] reported that activity autocorrelation timescales do not depend on the adaption time constant, whereas Muscinelli et al.[31] report a positive relationship. It is important to note, however, that the autocorrelation of neural activity in networks with adaptation does not follow a simple exponential decay. One possible explanation of the apparent disagreement in results could lie in the different methods used to measure timescales in each study, thus highlighting the need to understand how timescale estimation methods impact the conclusions we draw from empirical data and computational simulations. In the next section, we review how incorporating network mechanisms, in addition to cellular and synaptic mechanisms, could help to better explain observations of diverse and longer neural timescales.
@@ 3.2 Network and connectivity mechanisms
Neurons in the brain are embedded within rich network structures[63, 64] that affect the correlation patterns of neural dynamics[65, 66, 67, 68]. Therefore, a mechanistic understanding of neural timescales may require the context of interactions within a network. Indeed, computational models have shown that the strength and structure of network connectivity can influence the autocorrelation timescale of neural activity. Slower timescales arise from stronger[69, 46, 70, 71, 72] or more clustered (e.g., [73, 74]) connectivity (for a review see [75]). Strong recurrent connectivity in rate networks shifts the network’s dynamical state closer to a critical transition, increasing neural timescales [76, 77, 46, 32, 8]. In spiking networks, clustered connectivity creates metastable dynamics, and noise-induced transitions between these states give rise to slow timescales[73]. Metastable dynamics can also emerge in homeostatically regulated spiking networks[78]. Such structured connectivity leads to multiple dominant timescales in local dynamics, creating autocorrelation shapes beyond a single exponential decay[79, 8]. Each timescale is created by neural interactions on different spatial scales, such that slower timescales arise from interactions on a larger spatial scale, consistent with similar observations and theories on the spatiotemporal scale of oscillatory brain activity[80]. Introducing heterogeneity in connectivity strengths gives rise to distinct timescales in each network unit[47, 48]. In addition to recurrent interactions within one brain area, long-range connections between areas are essential for explaining the observed hierarchy of timescales[46].
@@
Recurrent inhibition also plays an important part in forming network-mediated neural timescales. Strong inhibitory-to-inhibitory connections allow for slow neural timescales required for working memory maintenance in a recurrent network model[81]. Moreover, in a detailed biophysical model of excitatory and inhibitory neurons, inhibition controls the stability and transitions of metastable states, giving rise to slow timescales[62]. In this model, the full shape of neural autocorrelations in frontal areas is explained by a combination of cellular (i.e., after-hyperpolarization potassium and inhibitory GABA-B conductance) and network mechanisms. In addition, recurrent inhibition together with membrane and synaptic time constants can account for distinct temporal dynamics in spiking activity and LFP[82].
Overall, computational modeling results have implicated a combination of synaptic, cellular, and network mechanisms to be partially responsible for the emergence of diverse neural timescales in the brain and across different modalities. While some of these mechanisms have found empirical support, either via testing different model predictions with experiments[46, 62, 8, 81] or through direct measurements and perturbations of the relevant physiological properties[54], and indirect and post-hoc correlations of neural timescales against, e.g., expression of ion channel and synaptic receptor genes (e.g.,[53, 11]), continued interaction between computational predictions and experimental measurements will further shed light on the mechanistic factors underlying variations in intrinsic neural timescales.
3.3 Modulatory mechanisms for dynamically shaping timescales
The mechanisms discussed so far explain how a specific range of timescales can emerge in neural dynamics due to “static” properties of single neurons and networks. However, neural timescales are functionally dynamic, and have been shown to adapt to task requirements[11, 17, 62, 49, 9, 51]. Here, we discuss how neuromodulatory processes, changes to feedforward and feedback inputs, and other processes can modulate cellular and network properties, and therefore the resulting activity timescales we measure.
@@ 3.3 Modulatory mechanisms for dynamically shaping timescales
Neuromodulators can impact the biophysical properties of individual neurons (e.g., the response time course) and their network interactions (e.g., gain modulation) (see [83, 84] for more extended reviews of the relation between the two and their impact on cortical computation), both of which can alter neural timescales. Synaptic time constants can be altered by neuromodulators such as acetylcholine (ACh)[85], which can also modify the efficacy of synaptic interactions across the brain in a selective manner[86, 87]. An increase in ACh strengthens the thalamocortical synaptic efficacy by affecting nicotinic receptors. At the same time, it weakens lateral cortical interactions while increasing the inhibitory drive by affecting muscarinic receptors. Similarly, the different contribution of glutamatergic receptors (e.g., NMDA and AMPA) in feedforward and recurrent pathways affects the time course of neural population responses in distinct ways[88].
External (e.g., sensory) inputs to a network can also modulate neural timescales through a variety of mechanisms. Input can bring individual neurons to a high-conductance state (lower resistance), which in turn reduces the membrane time constant[89] facilitating the distinction of distant synaptic inputs (for a review, see [90]). Moreover, in the presence of non-linear recurrent interactions, the input can modulate the global state of network dynamics[68] and alter activity timescales[8, 79, 49, 50]. The spatial and temporal scales (e.g., size and duration) of the stimulus can also distinctly impact fast and slow neural timescales. As a simple example, the activity of a network that tracks external signals over time will trivially exhibit the same timescales. Furthermore, the spatial extent of stimuli amplifies correlations at fast timescales while reducing them at slow timescales[91]. In spiking networks, populations with fast (slow) intrinsic timescales respond stronger to faster (slower) components of the input[48]. Thus, depending on the spatial or temporal profile of the input, different timescales might dominate the recorded neural activity.
The models discussed in this section suggest potential mechanisms underlying the diversity and flexibility of neural timescales. However, for these models to be more faithful representations of neural circuits, they need to reproduce not only the experimentally observed timescales but also capture other aspects of neural dynamics (e.g., realistic firing rates). Indeed, considering additional constraints can help to dissociate between various mechanisms that can explain the same range of neural timescales[8]. In addition, constraining computational models that were not specifically designed to explain neural timescales with realistic timescale measurements can provide new insights into different circuit mechanisms. For instance, constraining models to produce realistic timescales has been instrumental in developing models explaining the mechanisms underlying noise correlations observed in spiking activity[73, 68]. Given the relevance of neural timescales for brain computations and behavior, checking for realistic timescales (e.g., similar to realistic neural firing rates) in computational models can be used as a test for the plausibility of proposed models, and to contrast different mechanistic models which give rise to similar dynamical features, as we discuss in the next section.
4 Computational benefits of the diversity and flexibility of timescales
@@
As previously mentioned, recent experimental studies have consistently observed that timescales of both single-unit and population dynamics correlate with variables such as behavioral timescales and task performance, thus implicating neural timescales to be a signature of dynamical properties necessary for relevant computations [1, 2, 3]. However, to make causal claims about the necessity and sufficiency of dynamics operating on different timescales for network computation, perturbation experiments are required to assess the functional impact of timescales over and above, e.g., firing rate modulations, which can be challenging in vivo.
@@
Recently, task-performing artificial neural network (ANN) models offer a viable, complementary direction that circumvents the practical challenges of manipulating timescale-related variables in the in vivo brain, while enabling us to probe loss- or gain-of-function in experimental task paradigms. In computational and cognitive neuroscience, these networks are typically designed to mimic the input-output behavior observed in animal or human participants performing a task[92, 93]. To match the target output, the parameters of ANNs are either “hand-crafted” when possible (e.g., for simple tasks) or “trained” using an optimization algorithm minimizing a loss function (e.g., for larger networks and more complex tasks). The networks can then be further analyzed and perturbed to generate hypotheses for the mechanisms of computation at hand (as reviewed in, e.g., [94, 95]).
As a result, ANNs have been used to ask whether mechanisms shaping dynamics, such as connectivity, are important for computation, and whether they are preserved between artificial and biological networks performing the same task. Various types of ANNs—including spiking neural networks (SNN), rate-based recurrent neural networks (RNNs), and deep learning networks (DL)—can be leveraged to perform task-relevant computations while we explicitly measure, optimize for, and/or perturb their single-neuron and network timescales (Fig. 4a).
@@
In this section, we review two broad categories of timescale-related investigations using ANNs: First, we discuss those that study neural timescales as an emergent characteristic of task-performing networks, including both recent ones employing task-optimized networks, as well as classical works on dynamical properties of “hand-tuned” networks. Second, we discuss recent studies that explicitly optimize timescale-relevant parameters to enhance computational capacity in artificial networks, including biologically plausible spiking and rate networks, and more abstract models in the context of machine learning.
@@
Fig. 4: Task-performing models reveal functional relevance of timescales. a. Network models are set up to to achieve the input-output transformations similar to tasks performed by animals and humans. b. Certain geometry of population dynamics are desirable for network computation, such as fixed points, line attractors, and limit cycles. c. Model parameters, such as network weights and neuronal time constants, can be manually tuned when feasible or computationally optimized via gradient descent. d. Activity timescales and fitted time constants in network models match experimental data.

### v2

@@ 3 Mechanistic models of timescales in the brain
Mechanistic models, when constrained by experimental observations of neural timescales, can provide insights into how the diversity of timescales arises within and across different brain areas. Such models have so far been used to explain the following experimentally observed properties of neural timescales: (i) the hierarchical structure of neural timescales across different brain areas [68], (ii) the diversity of timescales within one brain area [69, 70, 24], and (iii) the flexibility of timescales to change according to different task variables [24, 9, 71, 72]. Mechanisms for capturing diverse intrinsic timescales are generally categorized into cellular and synaptic mechanisms on the level of individual neurons and network-mediated mechanisms through interactions between neurons (Fig. 4). At the same time, other processes, such as bottom-up (e.g., sensory) or top-down (e.g., from higher cortical areas) inputs and neuromodulatory mechanisms, can dynamically modulate timescales and give rise to stimulus- and task-dependent response timescales. Moreover, different models capture neural timescales on different spatial scales of observations, e.g., timescales of single neurons versus population activity. Thus, comparing these models can reveal how neural timescales across different scales and modalities can be related to each other.
@@ 3.1 Cellular and synaptic mechanisms
In computational models, timescales are often controlled by incorporating various biophysical time constants of cellular and synaptic processes in the dynamics of individual neurons that reflect heterogeneous properties of ionic channels and synapses [73, 74]. Most dynamical system models include a time constant in the dynamical equation of each neuron, interpreted as the membrane time constant ($5-50$ ms [75]) (Box 2). Slower timescales can be implemented by introducing other biophysical features such as slow synaptic time constants [76, 56, 77], timescales of short-term synaptic depression and facilitation [78, 79], or adaptation mechanisms [80, 81, 82, 56]. Spike-frequency adaptation, in particular, spans a wide range of fast and slow timescales [83], which could follow a power law [82]. Diverse adaptation timescales are suggested to lead to efficient coding in individual cortical neurons driven by behaviorally relevant signals [82] and facilitate computations on temporally dispersed inputs [84].
Box 2. Activity timescale as an estimated quantity vs. time constant as a parameter “Neural timescales” in the literature, for the most part, refer to the estimated autocorrelation decay timescale of measured neural activity, i.e., by describing the autocorrelation function of spike counts or LFP time series as $\textrm{ACF}(t)=e^{\frac{-t}{\tau_{\textrm{fit}}}}$. Methods discussed in Section 2 generally aim to find the value of $\tau_{\textrm{fit}}$ (or multiple $\tau_{\textrm{fit}}$s) that best fit the data. On the other hand, mechanistic (i.e. Hodgkin-Huxley) and functional (i.e., rate-based recurrent neural networks) models in Sections 3 and 4 typically include a time constant parameter that describes how the impact of input into a neuron decays over time, e.g., $\tau_{\textrm{param}}\frac{dV}{dt}=-(V(t)-V_{\textrm{rest}})+I(t)$, where $I(t)$ includes time-varying bottom-up (stimulus-driven), top-down, and recurrent inputs. It is important to note that the former, estimated $\tau_{\textrm{fit}}$ does not necessarily need to correspond to the latter, underlying model parameter $\tau_{\textrm{param}}$, and in general, does not. Section 3 outlines multiple mechanisms, including those unrelated to time constants, such as network connectivity, that contribute to shaping $\tau_{\textrm{fit}}$ we observe in data (see also Box 3).
@@
However, biophysical time constants at the single-neuron level cannot completely explain the temporal structure of neural dynamics. For instance, a Hodgkin-Huxley model of an isolated neuron with detailed dynamics of membrane potential and ionic and synaptic currents cannot capture the full autocorrelation shape of neural activity recorded in prefrontal areas [17]. Biophysical time constants alone define the autocorrelation timescale of neural activity—henceforth referred to as activity timescale (Box 2)—when a neuron is isolated, but when neurons are embedded in a network, their activity timescale may differ significantly from their biophysical time constants. At the same time, distinct biophysical mechanisms may interact differently in the presence of network interactions compared to isolated neurons. In particular, while activity timescales of network-embedded neurons increase with the synaptic time constant [56], their relationship to the adaptation time constant is less straightforward: Analyzing similar models of randomly connected excitatory and inhibitory units, Beiran and Ostojic [56] reported that activity autocorrelation timescales do not depend on the adaptation time constant, whereas Muscinelli et al. [57] report a positive relationship. It is important to note, however, that the autocorrelation of neural activity in networks with adaptation does not follow a simple exponential decay. One possible explanation of the apparent disagreement in results could lie in the different methods used to measure timescales in each study, thus highlighting the need to understand how timescale estimation methods impact the conclusions we draw from empirical data and computational simulations. In the next section, we review how incorporating network mechanisms, in addition to cellular and synaptic mechanisms, could help to better explain observations of diverse and longer neural timescales.
@@ 3.2 Network and connectivity mechanisms
Neurons in the brain are embedded within rich networks [85, 86] that affect the correlation patterns of neural dynamics [87, 88, 89, 90]. Therefore, a mechanistic understanding of neural timescales requires the context of interactions within a network. Most models rely on recurrent (excitatory or inhibitory) connectivity to explain how diverse neural timescales emerge, but as we discuss later, feedforward networks with rich temporal inputs can also give rise to diverse and hierarchical timescales [91].
Recurrent connectivity (within or across brain areas) can create positive or negative feedback that controls activity timescales. This can be done, for example, through creating attractor dynamics that give rise to persistent activity (e.g., line attractors) [92, 93] or shifting the network dynamical state closer to a critical transition [94, 95, 68, 39, 24, 96, 97, 98, 99, 100]. By analyzing linear recurrent networks (i.e., linear dynamical systems), we can define and compute the activity timescales of neurons analytically based on eigenvalues of the connectivity matrix (Box 3). However, in nonlinear networks, such timescales are not uniquely defined and can only be specified for a given (linearized) dynamical state, such as steady-state dynamics around fixed points [93, 101, 102, 34, 35].
The strength and structure of recurrent excitation influence the amount of positive feedback in a network and, consequently, the autocorrelation timescale of neural activity. In general, timescales increase with the strength of excitation [103, 68, 104, 105, 106, 94, 95, 68, 39, 24]. Moreover, heterogeneity in connection strengths can give rise to distinct timescales across network units [69, 70, 107, 100, 16]. For fixed connection strengths, timescales can additionally be affected by the connectivity structure. For instance, clustered connectivity creates metastable dynamics, and noise-induced transitions between these states give rise to slow timescales [108, 109, 110]. Metastable dynamics can also emerge in homeostatically regulated spiking networks [111].
Box 3. Activity timescales of a linear network As an illustrative example, we review how to compute the activity timescale of a linear recurrent network (i.e., linear dynamical system). Consider the system $\frac{dx(t)}{dt}=Ax(t)$ where $x(t)$ is the state vector, e.g., neural activity at time $t$, and $A$ is a square connectivity matrix. To acquire the timescales of the system dynamics, we solve for the eigenvalues of the matrix $A$, $\lambda_{i}$, through the characteristic equation, $A\textbf{u}_{i}=\lambda_{i}\textbf{u}_{i}$. In general, eigenvalues are complex and have the form $\lambda_{i}=\textrm{Re}(\lambda_{i})+j\textrm{Im}(\lambda_{i})$. For a stable decaying mode, i.e., non-exploding, the real part of the associated eigenvalue must be negative, and its decay timescale is defined as $\tau_{i}=-\frac{1}{\textrm{Re}(\lambda_{i})}.$ The dominant activity timescale of the entire system (i.e., estimated as $\tau_{\textrm{fit}}$) is then typically determined by the mode whose eigenvalue has the largest real part. Furthermore, for a linear network composed of units with fixed membrane time constant $\tau_{\textrm{param}}$, the system dynamics follow $\tau_{param}\frac{dx(t)}{dt}=Ax(t).$ Assuming $\tau_{\textrm{param}}$ to be the same for all neurons, the activity timescale is scaled accordingly, $\tau_{i}=-\frac{\tau_{\textrm{param}}}{\textrm{Re}(\lambda_{i})}.$ This approach has been applied to connectivity matrices incorporating further biological constraints, e.g., random networks obeying Dale’s law [219]. For nonlinear networks, it is necessary to first linearize the system around a specific point in state space before conducting the above analysis (see, e.g., [102]).
Furthermore, spatially structured connectivity leads to multiple dominant timescales in local dynamics, creating autocorrelation shapes beyond a single exponential decay [112, 24]. In these networks, each timescale is created by neural interactions on different spatial scales, such that slower timescales arise from interactions on a larger spatial scale, consistent with similar observations and theories on the spatiotemporal scale of oscillatory brain activity [113].
@@
Negative feedback via recurrent inhibition also plays an important part in forming neural timescales. Strong inhibitory-to-inhibitory connections allow for slow neural timescales required for working memory maintenance in a recurrent network model [114]. Moreover, in a detailed biophysical model of excitatory and inhibitory neurons, inhibition controls the stability and transitions of metastable states, giving rise to slow timescales [17]. In this model, the full shape of neural autocorrelations in frontal areas is explained by a combination of cellular (i.e., after-hyperpolarization potassium and inhibitory GABA-B conductance) and network mechanisms. In addition, recurrent inhibition together with membrane and synaptic time constants can account for distinct temporal dynamics in spiking activity and LFP [115].
3.3 Input-driven and modulatory mechanisms for dynamically shaping timescales
The mechanisms discussed so far explain how a specific range of timescales can emerge due to “static” properties of single neurons and networks. However, neural timescales are themselves dynamic, and can adapt to task demands [5, 23, 17, 9, 41, 72, 32, 71]. Here, we discuss how external inputs, including bottom-up sensory inputs and top-down inputs from higher brain areas, as well as neuromodulators such as acetylcholine (ACh) can alter neural timescales.
External input can have both direct and modulatory effects on the resulting timescales of a population. At the cellular level, incoming input can alter biophysical properties such as by bringing individual neurons to a high-conductance state (lower resistance), which reduces the membrane time constant [116] and facilitates the distinction of distant synaptic inputs [117]. The spatial and temporal statistics of the input itself can also impact response timescales measured from neural activity. Different features of sensory inputs can vary at different timescales (e.g., moving edges vs. persistent object identity). For example, theoretical work using feedforward models constructed based on slow feature analysis has suggested that fast input timescales are encoded in lower levels of visual hierarchy, while slower timescales are encoded in higher levels [91, 118, 119, 120, 121] (for a review, see [122]). In these models, learning invariant or slowly varying features of sensory stimuli gives rise to progressively slower representations and temporal stability of the stimulus. These model predictions have been confirmed experimentally in the rodent visual hierarchy [32]. Similar behavior can arise in recurrent networks, where populations with fast (slow) intrinsic timescales respond stronger to faster (slower) components of the input [70]. Furthermore, the spatial scale of stimuli modulates the timescale of population spiking activity by affecting feedforward and recurrent interactions [123]. Thus, depending on the spatiotemporal profile of the input, different timescales thereof may be encoded and reflected in neural activity.
Even without embedded temporal dynamics, input can modulate activity timescales by shifting the dynamical state of nonlinear networks. As discussed above, timescales can be estimated for nonlinear networks at a specific state. Different input levels thus place the network in different states, which can alter the “local” timescales [112, 93]. This mechanism could explain the modulation of neural timescales during selective-attention [24], aggression [34], and timing tasks [9, 71]. Moreover, it can capture how the organization of brain-wide response timescales during tasks deviate from the expected hierarchy of intrinsic timescales [93]. An interesting example is the sharp transition in response timescales from the middle temporal (MT) area to the medial superior temporal (MST) during working memory delay, despite their closeness in the visual hierarchy [8]. Large-scale cortical models suggest that this transition arises from nonlinear interactions between different brain areas, where distinct intrinsic and delay-activity response timescales are shaped by dynamics in different fixed points [93].
@@ 3.3 Input-driven and modulatory mechanisms for dynamically shaping timescales
Neuromodulatory projections from diffuse subcortical nuclei can also alter timescales as a result of their impact on cellular and network properties [124, 125, 126]. ACh, for instance, can change synaptic time constants and lead to persistent firing (e.g., during pain processing [127]). An increase in ACh, such as during selective attention, can also strengthen the thalamocortical synaptic efficacy by affecting nicotinic receptors, while weakening lateral cortical interactions by affecting muscarinic receptors [128]. Similarly, the different contribution of glutamatergic receptors in feedforward and recurrent pathways affects the time course of neural population responses in distinct ways [129]. Thus, neuromodulators can indirectly adapt neural timescales to task demands through selective gain modulation.
Altogether, a combination of cellular, network, input, and neuromodulatory mechanisms can shape the diversity and flexibility of neural timescales. We argue that constraining mechanistic computational models with realistic timescales can provide new insights into circuit mechanisms underlying neural population dynamics. While such models are typically challenging to constrain, recent open-source toolboxes leveraging approximate Bayesian computation (ABC) [23] and simulation-based inference (SBI) [60, 61] have enabled efficient discovery of data-consistent mechanistic models [26, 23, 130, 27, 28, 131]. Based on the available experimental measurements, one can constrain models with a single estimated timescale (e.g., average timescale of a brain area in Fig. 2a to average timescale of a network model), or with a distribution of timescales (e.g., diverse timescales within an area Fig. 2b to distribution of timescales within a model). As we discuss next, assessing for realistic timescales (e.g., similar to realistic firing rates) also allows us to test the plausibility of different models aimed to explain computation and function.
4 Timescale as signatures of dynamics supporting network computation
@@
Finally, we return to function: Recent experimental studies have consistently observed that timescales of both single-unit and population dynamics correlate with measurements such as behavioral timescales and task performance [1, 2, 3]. Neural timescales are thus implicated to be a signature of dynamical properties necessary for relevant computations. However, to make causal claims about the necessity and sufficiency of dynamics operating on different timescales for network computation, perturbation experiments are required to assess the functional impact of timescales over and above, e.g., firing rate modulations, which can be challenging in vivo.
@@
Recently, task-performing artificial neural network (ANN) models—including spiking and rate-based recurrent neural networks (SNNs and RNNs), as well as deep neural networks in machine learning—offer a viable, complementary direction: Such functional models circumvent the practical challenges of manipulating timescale-related variables in the in vivo brain, while enabling us to probe loss- or gain-of-function in experimental task paradigms. In computational and cognitive neuroscience, these networks are typically designed to mimic the input-output behavior observed in animal or human participants performing a task [132, 133] (Fig. 5a). To match the target output, the ANN parameters are either “hand-crafted” or analytically chosen when possible (e.g., for simple tasks; Fig. 5b, left), or “trained” using an optimization algorithm minimizing a loss function (e.g., for larger networks and more complex tasks; Fig. 5b, middle). Tuned networks can then be further analyzed and perturbed to generate hypotheses for the mechanisms of computation at hand, in particular with respect to properties of interest in phase space (Fig. 5c) [134, 135] and timescales of population dynamics (Fig. 5d).
@@
In this section, we review two broad categories of timescale-related investigations using ANNs: First, we discuss those that study neural timescales as an emergent characteristic of task-performing networks, including both classical works on dynamical properties of “hand-tuned” networks and recent ones employing task-optimized networks. Second, we discuss recent studies that explicitly optimize time constants and timescale-related parameters to enhance computational capacity in artificial networks (Fig. 5b, right), including biologically plausible SNNs and RNNs, as well as more powerful models in deep learning applications.
@@
Fig. 5: Task-performing models reveal functional relevance of timescales. a. Network models are set up to achieve input-output transformations similar to tasks performed by animals and humans. b. Model parameters, such as network weights and neuronal time constants, can be manually tuned when feasible or computationally optimized via gradient descent. c. A certain geometry of population dynamics is desirable for network computation, such as fixed points, line attractors, and limit cycles. d. Activity timescales and fitted time constants in network models match experimental data.